Start Me Up Inc. manufactures a caffeinated energy drink that sells for $4.30 each. The results for its first year of operations appear in the table below:

Projections
Number of drinks produced 49,000
Number of drinks sold 46,200
Direct materials per drink $ 0.58
Direct labor per drink $ 0.28
Variable manufacturing overhead per drink $ 0.18
Total fixed manufacturing overhead $ 30,380
Total fixed selling and administrative costs $ 51,500

Required:

1. Compute the operating income for the first year under full costing.

2. Compute the operating income for the first year under variable costing.

(For all requirements, do not round intermediate calculations.)

1. Operating income for the first year under full costing
2. Operating income for the first year under variable costing

PLEASE ANSWER ASAP, THANK YOU!
Answer #1
Full costing Variable costing
Direct materials 0.58 0.58
Direct labor 0.28 0.28
Variable manufacturing overhead 0.18 0.18
Fixed manufacturing overhead 0.62 0
Product cost per drink 1.66 1.04
1
Sales revenue 198660 =46200*4.30
Less: Cost of goods sold 76692 =46200*1.66
Less: Total fixed selling and administrative costs 51500
Operating income for the first year under full costing 70468
2
Sales revenue 198660 =46200*4.30
Less: Cost of goods sold 48048 =46200*1.04
Less: Total fixed selling and administrative costs 51500
Less: Total fixed manufacturing overhead 30380
Operating income for the first year under variable costing 68732